OpenStudy (anonymous):

how to solve 0.4x+6<=1.2x-5

OpenStudy (karatechopper):

x>=13.75

OpenStudy (mathteacher1729):

Add five to both sides \[0.4x + 11 \leq 1.2x\] Subtract \(0.4x\) from both sides. \[11 \leq 0.8x\] Divide by \(0.8\) on both sides \[12.75 \leq x\] Test this out to make sure it is correct. :)